A monochromatic beam of light of wavelength 500nm illuminates a double slit having a slit separation of 2.00mm. What is the angle of the second order bright fringe?


1
Expert's answer
2021-02-22T10:28:38-0500
"d\sin\theta=m\lambda\\\theta=\arctan{\frac{m\lambda}{d}}
\\\theta=\arctan{\frac{2(500)}{2\cdot10^6}}=0.03\degree"